Consider the following relative frequency density histogram. what is the area of all the shaded rectangles in the histogram? what proportion of the data is at most 0.4? what proportion of the data is greater than 0.8?
Answer #1

a) Area of all the shaded region = 0.2 * 1 + 0.2 * 1.5 + 0.4 * 0.75 + 0.2 * 1 = 1

b) P(at most 0.4) = P(X < 0.4) = 0.2 * 1 + 0.2 * 1.5 = 0.5

c) P(greater than 0.8) = P(X > 0.8) = 0.2 * 1 = 0.2